The probability that machine P will breakdown is 2/7 and the probability that machine Q will breakdown is 3/11.Find the probability that at least on machine will be working


1
Expert's answer
2021-11-05T03:05:19-0400

The probability that machine P will breakdown is "\frac{2}{7}"

The probability that machine P will not breakdown is "\frac{5}{7}"

The probability that machine Q will breakdown is "\frac{3}{11}"

The probability that machine Q will not breakdown is "\frac{8}{11}"

The probability that at least one machine will be working is "PQ'+P'Q+PQ= (\frac{2}{7}*\frac{8}{11})+(\frac{5}{7}*\frac{3}{11})+(\frac{2}{7}*\frac{3}{11})"

"=\frac{16}{77}+\left(\frac{5}{7}\cdot \frac{3}{11}\right)+\left(\frac{2}{7}\cdot \frac{3}{11}\right)\\
=\frac{16}{77}+\frac{15}{77}+\left(\frac{2}{7}\cdot \frac{3}{11}\right)\\
=\frac{16}{77}+\frac{15}{77}+\frac{6}{77}\\
=\frac{37}{77}"